Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2021-04-02

Observers:  Priscila Pessi, Ayan Mitra
Support: Phil Wiseman, Chris Frohmaier, Tomás Müller, Marcus Toy 
TAT: Takashi Nagao, Lluís Galbany, Shubham Srivastav

EFOSC2 Calibrations: 
Bias, dome_flats, arcs.
Skyflats but there were a lot of clouds so they might not be useful.

Weather:
Start: humidity 33% , wind  4.7 m/s, seeing 0.8 
Middle: humidity 29%, wind 6.3 m/s, seeing 0.6
End: humidity 30%, wind  10.2 m/s, seeing 0.9

Comments:
The telescope closed at 7:30UT for one hour due to clouds.
We opened again at 8:28UT

 UTTarget 
 TypeSetup SlitExposure (s)Airmass
Seeing 
Notes
23:44L745aSTANDARDg11+g13+g161.0/1.56X1001.00.9No image analysis because of clouds.
 23:54 FOCUS       
00:07AT2021blzFollow Upg131.027001.70.9 
01:04SN2020oczFollow Upg131.518001.61.7
01:42AT2021gnoFollow Upg11+gr161.52000s each1.71.4
 03:12 FOCUS      
03:15AT2021fxyFollow Upg11+gr161.0600s each1.11.2
04:03AT2021hkwClassificationg131.04501.41.2Very weak signal. Probably better to try again later.
04:20SDSS J1350+2916Follow Upgr11+gr161.0/1.52x1800s each2.20.6During the second integration of gr11 there might have been a cloud passing or something because the seeing jumped suddenly to ~2". After that the seeing improved but stayed at 1.6 so we used slit 1.5 for gr16
06:56gd153STANDARDg11+g13+g161.06x1002.01.0
       08:34
 AT2021elz classification gr13 1.0 1300 1.3 1.2 Very noisy
 09:07 AT2021hku classification gr13 1.0 500 1.4 0.9Ia
Comments